### v1

@@ III.4 Results
We first consider the first row of Fig.3, which shows the results for $\ell=0,m=0$. We see that the value of the estimated $\bar{B}_{0}^{0}$ is minimum for equilateral triangles $((\mu,t)=(0.5,1))$, and it increases as $\mu$ increases towards linear triangles $(\mu=1)$. The top-most curve of the right panel shows the analytical prediction $[\bar{B}_{0}^{0}]_{\rm d}(t)$ along the right boundary $(\mu\approx 1)$ of the $\mu-t$ plane. We see that $\bar{B}_{0}^{0}$ has the largest values along this curve. The value of $[\bar{B}_{0}^{0}]_{\rm d}(t)$ further increases as we move from stretched triangles $((\mu,t)=(1,0.5))$ to squeezed triangles $((\mu,t)=(1,1))$ along this curve. The bottom-most curve shows $[\bar{B}_{0}^{0}]_{\rm d}(\mu)$ along the top boundary $(t\approx 1)$ of the $\mu-t$ plane. Proceeding from equilateral to squeezed triangles along this curve, we see that $[\bar{B}_{0}^{0}]_{\rm d}$ increases gradually till $\mu\approx 0.9$, beyond which it increases sharply close to the squeezed limit. The middle curve shows $[\bar{B}_{0}^{0}]_{\rm d}(\mu)$ along the bottom boundary $(2\mu t\approx 1)$ of the $\mu-t$ plane. We see that the value of $[\bar{B}_{0}^{0}]_{\rm d}$ increases gradually as we proceed from equilateral to stretched triangles along this curve. In all cases, the estimated values $\bar{B}_{\ell}^{m}$ coincides with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$, and the fractional deviations are within the range $[-0.033,0.033]$. Considering the middle panel, we find that $\Delta^{0}_{0}\leq 3$ across most of the $\mu-t$ plane, which indicates that the differences between the estimated values and the analytical predictions are largely within the expected statistical fluctuations. We have a few bins where $3<\Delta^{0}_{0}\leq 5$, which also may be consistent with the expected statistical fluctuations. In addition, there are a few bins where $5<\Delta^{0}_{0}\leq 7$ where the deviations are in excess of those expected from statistical fluctuations. The $4$ bins where $5<\Delta^{0}_{0}\leq 7$ are all located near the squeezed limit, and a similar behaviour has been reported in earlier work [34]. The reason for this behaviour is that the bispectrum is very sensitive to the triangle shape near the squeezed limit, and the binned estimator includes triangles of different shapes due to the finite width of the annular rings (Fig. 1). In contrast, the analytical estimate only considers the bispectrum at the mean value of $(k_{1\perp},k_{2\perp},k_{3\perp},k_{1\parallel},k_{2\parallel})$ corresponding to the combination of three annular rings.
@@
We next consider the quadrupole moment ($\ell=2$) with $m=0$, which is shown in the the second row of Fig.3. We see that the values of $\bar{B}_{2}^{0}$ are comparable to those of $\bar{B}_{0}^{0}$. The results for $\bar{B}_{2}^{0}$ are also very similar to those for $\bar{B}_{0}^{0}$, and the right and left panels of the second row are very similar to the corresponding panels of the first row. Considering the middle panel, we see that the deviations from the analytical predictions are within the expected statistical fluctuations $(\Delta_{2}^{0}\leq 3)$ for nearly all the bins, with the exception of two bins (both at $\mu\approx 1$) where $3<\Delta^{0}_{0}\leq 5$.
@@
The third and fourth rows of Fig.3 show the quarupole moment ($\ell=2$) with $m=1$ and $m=2$ respectively. Unlike $\bar{B}_{0}^{0}$ and $\bar{B}_{2}^{0}$, which are positive throughout, we see that $\bar{B}_{2}^{1}$ and $\bar{B}_{2}^{2}$ have both positive and negative values. The magnitude of $\bar{B}_{2}^{1}$ and $\bar{B}_{2}^{2}$ are $\sim 4$ times smaller than those of $\bar{B}_{0}^{0}$ and $\bar{B}_{2}^{0}$. Further, the patterns in $\mu-t$ plane are also quite different. Considering $\bar{B}_{2}^{1}$, we see that it has small positive values for acute triangles ($\mu<t$), and negative values for obtuse triangles ($\mu>t$), with values close to zero for right-angled triangles $(\mu=t)$. Considering the right panel, we see that that magnitude is largest for squeezed triangles where $\bar{B}_{2}^{1}$ is negative. We also note that the values of $\bar{B}_{2}^{1}$ are in good agreement with those of $[\bar{B}_{2}^{1}]_{\rm d}$. The middle panel shows that $\Delta_{2}^{1}\leq 3$ for all the bins except the squeezed triangles where $3<\Delta_{2}^{1}\leq 5$. Considering $\bar{B}_{2}^{2}$, we see that the results are very similar to those for $\bar{B}_{2}^{1}$. However, there is the difference that the values of $\bar{B}_{2}^{2}$ are positive over a large region of the $\mu-t$ plane as compared to $\bar{B}_{2}^{1}$. We also see that $\Delta_{2}^{2}\leq 3$ for nearly all the bins except a few ($\sim 6$) where $3<\Delta_{2}^{2}\leq 5$.
@@
Overall, we see that the estimated values $\bar{B}_{\ell}^{m}$ are in good agreement with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$ for the monople moment $(\ell=0,m=0)$, and all the quadrupole moments $(\ell=2,m=0,1,2)$. We have carried out the same analysis for all the higher multipole moments, namely $(\ell=4,m=0,1,2,3,4)$ and $(\ell=6,m=0,1,2,3,4,5,6)$, that are predicted to be non-zero for our model [25]. The results for $\ell=4$ and $6$ are respectively presented in Figures 4 and 5 of Appendix A. In all the cases, we find that the values of $\bar{B}_{\ell}^{m}$ are in good agreement with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$. The values of $\Delta_{\ell}^{m}$ are mostly in the range $\Delta_{\ell}^{m}\leq 3$, and they are in the range $3<\Delta_{\ell}^{m}\leq 5$ for a few bins, indicating that the deviations between $\bar{B}_{\ell}^{m}$ and $[\bar{B}_{\ell}^{m}]_{\rm d}$ are consistent with that expected from statistical fluctuations.
@@ IV Summary and discussion
The bispectrum is the lowest-order statistic sensitive to non-Gaussianity in the underlying density field. This is particularly relevant for the late-time matter distribution where non-Gaussianity develops due to non-linear evolution [37], and the bispectrum contains a wealth of cosmological information [19, 10, 38, 21, 20, 30]. The redshifted 21-cm signal from the Epoch of Reionization is also predicted to be highly non-Gaussian due to the emergence of large ionized regions [39], and the bispectrum [40, 41, 42, 35] is expected to carry a plethora of information regarding the nature of the ionizing sources. It is therefore crucial to go beyond the power spectrum, and consider the bispectrum, and other higher-order statistics, to extract the maximum information from observational data.
@@
The bispectrum is a function of the closed triangles formed by three ${\bm{k}}$ vectors. Cosmological density fields are expected to be isotropic, and the bispectrum is predicted to depend only on the size and shape of the triangle, which we have parameterized using three parameters $k_{1}$ and $(\mu,t)$ respectively. However, due to redshift space distortion, the observed bispectrum also depends on how the triangle is oriented with respect to the LoS, and we require five parameters to parameterize the redshift space bispectrum. Here we have considered the redshift space bispectrum $B^{s}(\hat{\textbf{p}},k_{1},\mu,t)$, where the unit vector $\hat{\textbf{p}}$ [25] quantifies the orientation of the triangle with respect to the LoS. We have decomposed $B^{s}(\hat{\textbf{p}},k_{1},\mu,t)$ into the spherical harmonic basis $Y_{\ell}^{m}(\hat{\bm{p}})$ (eq. 4), and use the bispectrum multipole moments $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ to quantify the RSD anisotropy.
@@
It is relatively straightforward to compute the bispectrum by looping over the available ${\bm{k}}$ vectors, however, this is computationally very expensive as the computing time scales as $\sim N_{g}^{6}$ for a 3D data cube of $N_{g}$ grids in each dimension. This limitation is overcome by several FFT-based fast bispectrum estimators where the computational time scales as $\sim N_{g}^{3}\log{(N_{g})}$ ) [32, 33, 13]. However, these estimators consider spherical shells in ${\bm{k}}$ space, and they are restricted to the $m=0$ multipoles [43, 20] of the redshift space bispectrum. Considering the multipole moments with $\ell\leq 6$, we only have four non-zero multipole moments if we are restricted to $m=0$, whereas we have a total of sixteen non-zero multipole moments if we also include the ones with $m\neq 0$ [26]. The difficulty arises due to the $e^{im\phi}$ term in $Y_{\ell}^{m}(\hat{\bm{p}})$, as it is not possible to evaluate this with spherical shells in ${\bm{k}}$ space. We have overcome this by dividing the ${\bm{k}}$ space into $N_{g}$ planes perpendicular to the LoS, and dividing each plane into $N_{r}$ annular rings. The combination of three rings and three planes contains all the triangles of a fixed size, shape, and orientation to the LoS. We have used FFT to evaluate the bispectrum for such a combination, and we average this over different orientations to evaluate the bispectrum multipole moments (eq. 5). The computational time to estimate the bispectrum for each combination of planes scales as $\sim N_{g}^{2}\log{(N_{g})}$ ), and we have $\sim N_{g}^{3}\log{(N_{g})}$ ) considering all $N_{g}$ planes.
@@
We have validated our estimator using simulations. We have considered 50 realizations of a mildly non-Gaussian field that incorporates redshift space distortion. The analytical expressions for bispectrum multipoles of this field are presented in [25]. We have used the estimator to calculate all the bispectrum multipoles that are predicted to be non-zero (up to $\ell=6,m=6$ in our case). We find that the estimated values $\bar{B}_{\ell}^{m}$ are in good agreement with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$ for all the cases. The deviations of the estimates from the analytical predictions are within ±$\pm 3\sigma$ for most of the bins, consistent with their arising from statistical fluctuations. In a few bins, the deviations are in the range ±$\pm 3\sigma$ to ±$\pm 5\sigma$, which also is possibly due to statistical fluctuations. In some rare cases, mostly near the squeezed limit, we have deviations in the range ±$\pm 5\sigma$ to ±$\pm 7\sigma$. However, in these cases, the fractions deviations are of the order of $1\%$ or smaller. In conclusion, our estimator is able to calculate all the non-zero bispectrum multipole moments at a high level of accuracy. It is anticipated that including the $m\neq 0$ multipoles will add to the information already available from galaxy redshift surveys, and that expected from future redshifted 21-cm observations.
@@ Acknowledgements
SSG acknowledges the support of the Prime Minister Research Fellowship (PMRF). We thank Abinash Kumar Shaw for the useful discussions. We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Shakti’ at IIT Kharagpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India.
@@
The simulated data and package involved in this work will be shared on reasonable request to the authors.
@@ Appendix A Results for $\ell=4$ and $6$
Fig.4 and 5 shows the results for $\ell=4$ and $6$ bispectrum multipoles respectively. The estimates for all the cases are in good statistical agreement with the analytical predictions. The statistical deviation is $\Delta_{\ell}^{m}\leq 3$ for most of the cases, and it is in the range $3<\Delta_{\ell}^{m}\leq 5$ for a few bins near linear triangles.

### v2

@@ III.4 Results
We first consider the first row of Fig. 4, which shows the results for $\ell=0,m=0$. We see that the value of the estimated $\bar{B}_{0}^{0}$ is minimum for equilateral triangles $((\mu,t)=(0.5,1))$, and it increases as $\mu$ increases towards linear triangles $(\mu=1)$. The top-most curve of the right panel shows the analytical prediction $[\bar{B}_{0}^{0}]_{\rm d}(t)$ along the right boundary $(\mu\approx 1)$ of the $\mu-t$ plane. We see that $\bar{B}_{0}^{0}$ has the largest values along this curve. The value of $[\bar{B}_{0}^{0}]_{\rm d}(t)$ further increases as we move from stretched triangles $((\mu,t)=(1,0.5))$ to squeezed triangles $((\mu,t)=(1,1))$ along this curve. The bottom-most curve shows $[\bar{B}_{0}^{0}]_{\rm d}(\mu)$ along the top boundary $(t\approx 1)$ of the $\mu-t$ plane. Proceeding from equilateral to squeezed triangles along this curve, we see that $[\bar{B}_{0}^{0}]_{\rm d}$ increases gradually until $\mu\approx 0.9$, beyond which it increases sharply close to the squeezed limit. The middle curve shows $[\bar{B}_{0}^{0}]_{\rm d}(\mu)$ along the bottom boundary $(2\mu t\approx 1)$ of the $\mu-t$ plane. We see that the value of $[\bar{B}_{0}^{0}]_{\rm d}$ increases gradually as we proceed from equilateral to stretched triangles along this curve. In all cases, the estimated values $\bar{B}_{\ell}^{m}$ coincide with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$, and the fractional deviations are within the range $[-0.033,0.033]$. Considering the middle panel, we find that $\Delta^{0}_{0}\leq 3$ across most of the $\mu-t$ plane, which indicates that the differences between the estimated values and the analytical predictions are largely within the expected statistical fluctuations. We have a few bins where $3<\Delta^{0}_{0}\leq 5$, which also may be consistent with the expected statistical fluctuations. In addition, there are a few bins where $5<\Delta^{0}_{0}\leq 7$ where the deviations are in excess of those expected from statistical fluctuations. The $3$ bins where $5<\Delta^{0}_{0}\leq 7$ are all located near the squeezed limit, and a similar behavior has been reported in earlier work [35]. The reason for this behavior is that the bispectrum is very sensitive to the triangle shape near the squeezed limit. For a particular combination of rings $(r_{1},r_{2},r_{3})$ and planes $(n_{1},n_{2},n_{3})$, the binned estimator (Eq. 8) includes triangles of different shapes due to the finite width of the annular rings (Fig. 1). In contrast, as discussed in Section III.3, the discrete analytic prediction $[\bar{B}_{\ell}^{m}]_{\rm d}$ only considers the bispectrum at the mean value of $(k_{\perp r_{1}},k_{\perp r_{2}},k_{\perp r_{3}},k_{\parallel n_{1}},k_{\parallel n_{2}})$ corresponding to the combination of three annular rings and planes. Note, however, that $[\bar{B}_{\ell}^{m}]_{\rm d}$ incorporates the second level of binning (Eq. 13) that is used to calculate $\hat{\bar{B}}_{\ell}^{m}$.
@@
We next consider the quadrupole moment ($\ell=2$) with $m=0$, which is shown in the second row of Fig. 4. We see that the values of $\bar{B}_{2}^{0}$ are comparable to those of $\bar{B}_{0}^{0}$. The results for $\bar{B}_{2}^{0}$ are also very similar to those for $\bar{B}_{0}^{0}$, and the right and left panels of the second row are very similar to the corresponding panels of the first row. Considering the middle panel, we see that the deviations from the analytical predictions are within the expected statistical fluctuations $(\Delta_{2}^{0}\leq 3)$ for nearly all the bins except the squeezed triangles where $3<\Delta^{0}_{0}\leq 5$.
@@
The third and fourth rows of Fig. 4 show the quadrupole moment ($\ell=2$) with $m=1$ and $m=2$, respectively. Unlike $\bar{B}_{0}^{0}$ and $\bar{B}_{2}^{0}$, which are positive throughout, we see that $\bar{B}_{2}^{1}$ and $\bar{B}_{2}^{2}$ have both positive and negative values. The magnitude of $\bar{B}_{2}^{1}$ and $\bar{B}_{2}^{2}$ are $\sim 4$ times smaller than those of $\bar{B}_{0}^{0}$ and $\bar{B}_{2}^{0}$. Further, the patterns in the $\mu-t$ plane are also quite different. Considering $\bar{B}_{2}^{1}$, we see that it has small positive values for acute triangles ($\mu<t$) and negative values for obtuse triangles ($\mu>t$), with values close to zero for right-angled triangles $(\mu=t)$. Considering the right panel, we see that magnitude is the largest for squeezed triangles where $\bar{B}_{2}^{1}$ is negative. We also note that the values of $\bar{B}_{2}^{1}$ are in good agreement with those of $[\bar{B}_{2}^{1}]_{\rm d}$. The middle panel shows that $\Delta_{2}^{1}\leq 3$ for all the bins, with the exception of two bins (both at $\mu\approx 1$) where $3<\Delta_{2}^{1}\leq 5$. Considering $\bar{B}_{2}^{2}$, we see that the results are very similar to those for $\bar{B}_{2}^{1}$. However, there is a difference that the values of $\bar{B}_{2}^{2}$ are positive over a large region of the $\mu-t$ plane as compared to $\bar{B}_{2}^{1}$. We also see that $\Delta_{2}^{2}\leq 3$ for nearly all the bins except a few ($\sim 5$) where $3<\Delta_{2}^{2}\leq 5$.
@@
Overall, we see that the estimated values $\bar{B}_{\ell}^{m}$ are in good agreement with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$ for the monopole moment $(\ell=0,m=0)$, and all the quadrupole moments $(\ell=2,m=0,1,2)$. We have carried out the same analysis for all the higher multipole moments, namely $(\ell=4,m=0,1,2,3,4)$ and $(\ell=6,m=0,1,2,3,4,5,6)$, that are predicted to be non-zero for our model [25]. The results for $\ell=4$ and $6$ are respectively presented in Figures 5 and 6 of Appendix A. In all the cases, we find that the values of $\bar{B}_{\ell}^{m}$ are in good agreement with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$. The values of $\Delta_{\ell}^{m}$ are mostly in the range $\Delta_{\ell}^{m}\leq 3$, and they are in the range $3<\Delta_{\ell}^{m}\leq 5$ for a few bins, indicating that the deviations between $\bar{B}_{\ell}^{m}$ and $[\bar{B}_{\ell}^{m}]_{\rm d}$ are consistent with that expected from statistical fluctuations.
@@ IV Summary and discussion
The bispectrum is the lowest-order statistic sensitive to non-Gaussianity in the underlying density field. This is particularly relevant for the late-time matter distribution where non-Gaussianity develops due to non-linear evolution [38], and the bispectrum contains a wealth of cosmological information [19, 10, 39, 29, 21, 20, 30]. The redshifted 21-cm signal from the Epoch of Reionization is also predicted to be highly non-Gaussian due to the emergence of large ionized regions [40], and the bispectrum [41, 42, 43, 36] is expected to carry a plethora of information regarding the nature of the ionizing sources. It is, therefore, crucial to go beyond the power spectrum and consider the bispectrum and other higher-order statistics to extract the maximum information from observational data.
@@
The bispectrum is a function of the closed triangles formed by three ${\bm{k}}$ vectors. Cosmological density fields are expected to be isotropic, and the bispectrum is predicted to depend only on the size and shape of the triangle, which we have parameterized using three parameters $k_{1}$ and $(\mu,t)$ respectively. However, due to redshift space distortion, the observed bispectrum also depends on how the triangle is oriented with respect to the LoS, and we require five parameters to parameterize the redshift space bispectrum. Here we have considered the redshift space bispectrum $B^{s}(\hat{\textbf{p}},k_{1},\mu,t)$, where the unit vector $\hat{\textbf{p}}$ [25] quantifies the orientation of the triangle with respect to the LoS. We have decomposed $B^{s}(\hat{\textbf{p}},k_{1},\mu,t)$ into the spherical harmonic basis $Y_{\ell}^{m}(\hat{\bm{p}})$ (Eq. 4), and used the bispectrum multipole moments $\bar{B}_{\ell}^{m}(k_{1},\mu,t)$ to quantify the RSD anisotropy.
@@
It is relatively straightforward to compute the bispectrum by looping over the available ${\bm{k}}$ vectors, however, this is computationally very expensive as the computing time scales as $\sim N_{g}^{6}$ for a 3D data cube of $N_{g}$ grids in each dimension. This limitation is overcome by several FFT-based fast bispectrum estimators where the computation time scales as $\sim N_{g}^{3}\log{(N_{g})}$ ) [33, 34, 13]. However, these estimators consider spherical shells in ${\bm{k}}$ space, and they are restricted to the $m=0$ multipoles [44, 20] of the redshift space bispectrum. Considering the multipole moments with $\ell\leq 6$, we only have four non-zero multipole moments if we are restricted to $m=0$, whereas we have a total of sixteen non-zero multipole moments if we also include the ones with $m\neq 0$ [26]. The difficulty arises due to the $e^{im\phi}$ term in $Y_{\ell}^{m}(\hat{\bm{p}})$, as it is not possible to evaluate this with spherical shells in ${\bm{k}}$ space. We have overcome this by dividing the ${\bm{k}}$ space into $N_{g}$ planes perpendicular to the LoS and dividing each plane into $N_{r}$ annular rings. The combination of three rings and three planes contains all the triangles of a fixed size, shape, and orientation with respect to the LoS. We have used the FFT to evaluate the bispectrum for such a combination, and we average this over different orientations to evaluate the bispectrum multipole moments (Eq. 5). The computation time for performing the FFT operations scale as $\mathcal{O}(N_{g}^{3})$, while estimating all the bispectrum multipoles for all possible closed triangle configurations scales as $\mathcal{O}(N_{g}^{4})$. Note that the estimator presented in this work is based on the plane-parallel approximation and assumes a fixed LoS. However, in realistic modern galaxy surveys that cover large angles of the sky, the LoS varies with the positions of galaxies, leading to the radial RSD. The effect due to the radial nature of the distortions may be important and can significantly influence the galaxy statistics. Yamamoto et al. [45] introduced a formalism to account for the varying LoS in the power spectrum multipoles, where the LoS is defined along the direction of one of the galaxies in a pair. Building on this, Scoccimarro [13] proposed an estimator for the $m=0$ multipoles of the bispectrum that incorporates these radial distortions. We plan to incorporate it in our estimator in the future.
@@
We have validated our estimator using simulations. We have considered 50 realizations of a mildly non-Gaussian field that incorporates redshift space distortion. The analytical expressions for the bispectrum multipoles of this field are presented in [25]. We have used the estimator to calculate all the bispectrum multipoles that are predicted to be non-zero (up to $\ell=6,m=6$ in our case). We find that the estimated values $\bar{B}_{\ell}^{m}$ are in good agreement with the analytical predictions $[\bar{B}_{\ell}^{m}]_{\rm d}$ for all the cases. The deviations of the estimates from the analytical predictions are within ±$\pm 3\sigma$ for most of the bins, consistent with their arising from statistical fluctuations. In a few bins, the deviations are in the range ±$\pm 3\sigma$ to ±$\pm 5\sigma$, which is also possibly due to statistical fluctuations. In some rare cases, mostly near the squeezed limit, we have deviations in the range ±$\pm 5\sigma$ to ±$\pm 7\sigma$. However, in these cases, the fractional deviations are still of the order of $1\%$ or smaller. In conclusion, our estimator is able to calculate all the non-zero bispectrum multipole moments at a high level of accuracy. It is anticipated that including the $m\neq 0$ multipoles will add to the information already available from galaxy redshift surveys, and that is expected from future redshifted 21-cm observations.
@@ Acknowledgements
The authors thank the anonymous reviewer for the valuable suggestions and comments. We thank Abinash Kumar Shaw for the useful discussions. SSG acknowledges the support of the Prime Minister Research Fellowship (PMRF). We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Shakti’ at IIT Kharagpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India.
@@
The simulated data and the package involved in this work will be shared on reasonable request to the corresponding authors.
@@ Appendix A Results for $\ell=4$ and $6$
Fig. 5 and 6 shows the results for $\ell=4$ and $6$ bispectrum multipoles respectively. Note that these multipole moments have a smaller SNR as compared to the $\ell=0$ and $2$ multipoles, and we have divided the $\mu-t$ plane into wider bins to account for this. While the interval $0.5\leq\mu,t\leq 1$ was divided into $12$ bins along each direction for both $\ell=0$ and $2$, we have used $8$ and $6$ bins for $\ell=4$ and $6$ respectively. For all the cases shown here, the estimates are in good statistical agreement with the analytical predictions. We find $\Delta_{\ell}^{m}\leq 3$ for most of the cases, and it is in the range $3<\Delta_{\ell}^{m}\leq 5$ for a few bins near linear triangles. These deviations are roughly consistent with those expected from statistical fluctuations.